OpenStudy (anonymous):

xy=2 Is this? a)symmetry at x-axis,y-axis, and/or at the origin

OpenStudy (anonymous):

rewrite as y = 2/x... recognize this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

also if f(-x) = f(x) then it is symmetric about y. if f(-x) = -f(x) it is symmetric about the origin. if f(x) = y or -y then it is symmetric about x.

OpenStudy (anonymous):

not or, and.

OpenStudy (anonymous):

that is, x: -> y and x -> -y then it is not a function but is symmetric about the x-axis.

OpenStudy (anonymous):

trying to keep up but im sure y=2/x is a function and i wanted to know if any of the combinations of symmetry x-axis,y-axis,origin(or just at least one of these) are possible

OpenStudy (anonymous):

learn to do this: \[f(x) = \frac{ 2 }{ x }\Rightarrow f(-x) = \frac{ 2 }{ -x }=-\frac{ 2 }{ x }\Rightarrow f(-x) = -f(x)\] so it's symmetric about the origin. since it's a function. it can't be symmetric about the x-axis.